OpenStudy (anonymous):

given that x is a normal variable with mean μ=47 and standard deviation σ=6.2, find P(x≤60), P(x≥50), P(50≤x≤60)

OpenStudy (zarkon):

\[P(X≤60)=P\left(\frac{X-47}{6.2}\leq\frac{60-47}{6.2}\right)=P(Z\leq2.0967741935484)=0.98199327897418\]

OpenStudy (anonymous):

is this for all three?

OpenStudy (zarkon):

no...that is just the first one

OpenStudy (anonymous):

oh ok